OpenStudy (anonymous):

Giving medal to whoever can help? To open these doors, you must match the number and type of solutions for the following two functions in standard form. f(x) = x2 + 6x – 16 g(x) = x2 +6x + 1 Match the following descriptions of the solutions to each of the functions above. Hint: they each have their own match. Two real irrationals solutions Two real rationals solutions After matching these functions, explain to Professor McMerlock how you know these functions meet each condition. Remember, he is a professor, so use complete sentences.

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

please...

OpenStudy (anonymous):

if discriminant D=b^2-4ac>0 and a perfect square,then roots are real ,rational and different. if D>0 and not a perfect square then roots are real and irrational.

OpenStudy (anonymous):

woah i didn't quite understand that..

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

for first D=b^2-4ac=6^2-4(1)(-16=36+64=100=10^2 what can you say now?

OpenStudy (anonymous):

correction d=6^2-4(1)(-16)=36+64=100=10^2

OpenStudy (anonymous):

I'm sorry i thought you were suppose to just match the two top ones with the bottom and then explain why each one is which..

OpenStudy (anonymous):

i have solved for f(x)

OpenStudy (anonymous):

oh ok so the answer for the f(x) is d=6^2-4(1)(-16)=36+64=100=10^2

OpenStudy (anonymous):

explanation i have given above ,you have to see. you see it is positive and a perfect square.

OpenStudy (anonymous):

what about for g(x)?

OpenStudy (anonymous):

sorry I'm bothering, ill give you the medal just please help with that last part..

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

?

OpenStudy (anonymous):

can someone at least just explain the bottom half for me?